Reactions of Laser Ablated Be Atoms with H_{2}O: Infrared Spectra of HOBeOH and H_{2}O-BeO in Solid Argon.

\underline{C. Thompson},^{a} and L. Andrews, University of Virginia, Dept. of Chemistry, McCormick Road, Charlottesville, VA 22901.

Products formed by reacting laser ablated Be atoms with H_{2}O included beryllium oxides seen in earlier Be + O_{2} experiments^{b} as well as novel molecules mentioned above. Similar products arise from the formation of BeO in both systems. Once formed BeO may either react with H_{2}O to form HOBeOH or form the H_{2}O-BeO complex. Products were identified by several modes in excellent agreement with DFT calculations and provide interesting comparisons to beryllium oxides such as OBeO and Ar-BeO.^{c} HOBeOH is calculated to have an HOOH dihedral angle of 86^{∘} comparable to H_{2}O_{2}. The molecules also has a linear OBeO subunit whose antisymmetric stretch is observed about 81 cm^{−1} to the blue of OBeO. Spectra for both species will be presented along with a brief comparison to the oxygen system. aCurrently a National Research Council-Naval Research Laboratory Associate. Naval Research Laboratory. Code 6110 4555 Overlook Avenue SW Washington DC 20375-5342. bThompson, C. A.; Andrews, L. J. Chem. Phys. \underline{100}, 8689 (1994). cThompson, C. A.; Andrews, L. J. Am. Chem. Soc. \underline{116}, 423 (1994).